# Robust covariance estimation and explainable outlier detection for matrix-valued data

### v1

@@ Abstract
The minimum covariance determinant (MCD) estimator is a popular method for robustly estimating the mean and covariance of multivariate data. We extend the MCD to the setting where the observations are matrices rather than vectors and introduce the matrix minimum covariance determinant (MMCD) estimators for robust parameter estimation. These estimators hold equivariance properties, achieve a high breakdown point, and are consistent under elliptical matrix-variate distributions. We have also developed an efficient algorithm with convergence guarantees to compute the MMCD estimators. Using the MMCD estimators, we can compute robust Mahalanobis distances that can be used for outlier detection. Those distances can be decomposed into outlyingness contributions from each cell, row, or column of a matrix-variate observation using Shapley values, a concept for outlier explanation recently introduced in the multivariate setting. Simulations and examples reveal the excellent properties and usefulness of the robust estimators.
Keywords Covariance with Kronecker structure  $\cdot$ Explainable artificial intelligence  $\cdot$ Image data  $\cdot$ Minimum covariance determinant  $\cdot$ Shapley values
@@ 1 Introduction
Thanks to modern data collection tools, the amount and complexity of available information are increasing rapidly, and often, matrix-valued data are observed. Compared to classical multivariate observations, where values for $p$ variables are recorded for one subject, matrix-valued observations are recorded on a grid of $p\times q$ variables. These are then naturally represented as a matrix with $p$ rows and $q$ columns. Some examples include image data, where $p$ and $q$ are given by the resolution of the image, or multivariate data measured on $p$ variables, where the measurements for a subject are available for $q$ replications (e.g., different time points, different spatial locations, different experimental conditions, etc.). Frequently, matrix-valued data are analyzed as classical multivariate data by stacking the matrix columns (or rows) to a vector of length $p\cdot q$. Thus, if $n$ observations are available, the data are arranged in a matrix of dimension $n\times pq$. Depending on the dimensions, this can create high-dimensional data, possibly with a sample size lower than the resulting dimensionality, which constitutes a limitation for several multivariate statistical methods.
@@
As an alternative to vectorizing matrix-valued observations, we model them under the assumption that they originate from a certain matrix-variate distribution. As in the multivariate case, elliptical matrix-variate distributions are studied frequently (Gupta and Varga,, 2012), with particular focus being placed on the most prominent representative, i.e., the matrix normal distribution (Dawid,, 1981). Formally, a random matrix $\bm{X}\in\mathbb{R}^{p\times q}$ follows a matrix normal distribution $\mathcal{MN}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$ with mean $\bm{\operatorname{M}}\in\mathbb{R}^{p\times q}$, row covariance $\bm{\Sigma}^{\operatorname{row}}\in\operatorname{PSD}(p)$, and column covariance $\bm{\Sigma}^{\operatorname{col}}\in\operatorname{PSD}(q)$, if and only if its vectorized version $\operatorname{vec}(\bm{X})\in\mathbb{R}^{pq}$ follows a multivariate normal distribution $\mathcal{N}(\operatorname{vec}(\bm{\operatorname{M}}),\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}})$ (Gupta and Nagar,, 1999). Here, $\operatorname{vec}(\cdot)$ is the vectorization operator, stacking the columns of a matrix on top of each other, $\otimes$ is the Kronecker product, and $\operatorname{PSD}(a)$, with $a\in\mathbb{N}$, denotes the class of all positive definite symmetric $a\times a$ matrices. The probability density function (pdf) of a matrix normal random variable $\bm{X}$ is given by
$f(\bm{X}|\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})=\frac{\exp(-\frac{1}{2}\operatorname{tr}(\bm{\Omega}^{\operatorname{col}}(\bm{X}-\bm{\operatorname{M}})^{\prime}\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}})))}{(2\pi)^{\nicefrac{{pq}}{{2}}}\det(\bm{\Sigma}^{\operatorname{col}})^{\nicefrac{{p}}{{2}}}\det(\bm{\Sigma}^{\operatorname{row}})^{\nicefrac{{q}}{{2}}}},$ (1)
with $\bm{\Omega}^{\operatorname{row}}=(\bm{\Sigma}^{\operatorname{row}})^{-1}$ and $\bm{\Omega}^{\operatorname{col}}=(\bm{\Sigma}^{\operatorname{col}})^{-1}$ denoting the precision matrices among the rows and columns, respectively.
@@
Regarding the estimation of location and covariance for an i.i.d. sample $\bm{\mathfrak{X}}=(\bm{X}_{1},\ldots,\bm{X}_{n})\in\mathbb{R}^{n\times p\times q}$, with $\bm{X}_{i}\sim\mathcal{MN}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$, we can either work with the vectorized observations or directly with the matrices. In the former setting, the existence and uniqueness of the maximum likelihood estimator (MLE) for the covariance is guaranteed almost surely if $n\geq pq+1$. However, this approach does not take advantage of the Kronecker structure of the covariance matrix and instead directly estimates the entire $pq$-dimensional matrix $\bm{\Sigma}$. In contrast, if we utilize the knowledge of the inherent data structure, we only need to estimate the $p$-dimensional rowwise covariance matrix $\bm{\Sigma}^{\operatorname{row}}$ and the $q$-dimensional columnwise covariance matrix $\bm{\Sigma}^{\operatorname{col}}$. For the matrix-variate sample $\bm{\mathfrak{X}}$, the MLEs for the mean, as well as for the rowwise and columnwise covariance are given by (Dutilleul,, 1999):
@@
Soloveychik and Trushin, (2016) showed that for $n$ i.i.d. samples from a continuous $p\times q$ matrix-variate distribution, there exists no unique maximum of the matrix normal likelihood function if $n<\max(\nicefrac{{p}}{{q}},\nicefrac{{q}}{{p}})+1$, and that a unique maximum exists almost surely if $n\geq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$. While there exists no closed-form solution for the MLEs of $\bm{\Sigma}^{\operatorname{row}}$ and $\bm{\Sigma}^{\operatorname{col}}$, an iterative estimation procedure was proposed by Dutilleul, (1999). The idea of the so-called flip-flop algorithm is to alternate between the computation of $\hat{\bm{\Sigma}}^{\operatorname{row}}$ and $\hat{\bm{\Sigma}}^{\operatorname{col}}$ based on Equations (3) and (4), respectively, until a convergence criterion is met. The algorithm is constructed such that positive definite estimates of subsequent iterations are nondecreasing in likelihood (Lu and Zimmerman,, 2005) and it converges almost surely to the unique maximum from any symmetric positive definite initialization of either $\hat{\bm{\Sigma}}^{\operatorname{row}}$ or $\hat{\bm{\Sigma}}^{\operatorname{col}}$, if $n\geq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$ (Soloveychik and Trushin,, 2016).
@@
In this paper, we are interested in a robust estimation of the parameters $\bm{\operatorname{M}}$, $\bm{\Sigma}^{\operatorname{row}}$, and $\bm{\Sigma}^{\operatorname{col}}$ of the matrix normal distribution. Existing proposals include a generalization of Tyler’s M-estimator (Tyler,, 1987) introduced by Soloveychik and Trushin, (2016), a robust estimator for structured covariance matrices with Kronecker structure as a particular case (Sun et al.,, 2016), distribution-free robust covariance estimation (Zhang et al.,, 2022), and ML estimation for the matrix t-distribution (Thompson et al.,, 2020).
We propose a new robust estimation method, the matrix-variate minimum covariance determinant (MMCD) estimator. This is an extension of the minimum covariance determinant (MCD) estimator (Rousseeuw,, 1985), a popular robust estimator of location and covariance for the case of multivariate data. We will show that the MMCD is equivariant under linear matrix transformations, consistent under elliptical distributions, and achieves a high breakdown point. It also achieves a higher breakdown point than the MCD when applied to vectorized matrix-valued observations. Furthermore, an algorithm similar to the Fast-MCD algorithm (Rousseeuw and Driessen,, 1999) is developed for the computation of the MMCD, which makes the methodology also accessible to higher-dimensional matrices.
@@
The robust MMCD estimators can then be employed for outlier detection using the Mahalanobis distances (Mahalanobis,, 1936) for matrix-valued observations. Because it is essential to understand the reasons for the outlyingness, we will extend the concept of Shapley values introduced in Mayrhofer and Filzmoser, (2023) for outlier explanation in the multivariate case to the matrix-variate setting. Shapley values (Shapley,, 1953) are well-known from explainable AI (Lundberg and Lee,, 2017), but their computation is usually time-consuming. Our proposal is computationally efficient, and the resulting Shapley values preserve their attractive properties (Shapley,, 1953).
@@
The paper is organized as follows. In Section 2, we introduce the MMCD estimators, then proceed to derive their theoretical properties in Section 3. Section 4 is devoted to computational details for the MMCD estimators. In Section 5, we propose Shapley values for outlier explanation and present their properties. In Sections 6 and 7, we illustrate the performance of the proposed methods on numerical simulations and real-world examples. Section 8 concludes our findings. The Supplement contains more theoretical background information in this context, proofs, technical derivations, and additional numerical results.
@@ 2 The MMCD estimators
The goal is to derive robust estimators for the parameters $\bm{\operatorname{M}}$, $\bm{\Sigma}^{\operatorname{row}}$, and $\bm{\Sigma}^{\operatorname{col}}$ in Equation (1) mimicking the approach of the MCD estimator, whose objective is to identify the subset of $h$ out of $n$ samples, with $\nicefrac{{n}}{{2}}\leq h\leq n$, whose sample covariance matrix has the smallest determinant. For the MMCD estimators, we use the matrix MLE estimators instead of the sample covariance of the vectorized data. Therefore, we use the connection between the MCD estimator and the MLEs for mean and covariance of the multivariate normal distribution, as pointed out in Raymaekers and Rousseeuw, (2023): Minimizing the determinant of the sample covariance matrix of a data subset is equivalent to minimizing the (log-)likelihood function for that subset. The latter corresponds to minimizing a trimmed (log-)likelihood, where trimming is done by introducing binary weights for the observations. This idea has been similarly used in the context of other robust estimators (e.g. Neykov et al.,, 2007; García-Escudero et al.,, 2010; Kurnaz et al.,, 2018). More details regarding the matrix-variate setting and the MCD procedure are provided in Supplement A.
The MLEs given in Equations (2)-(4) are not robust against outlying (matrix-variate) observations. Thus, we introduce weights $\bm{w}=(w_{1},\dots,w_{n})\in\mathbb{R}^{n}$ for a given sample $\bm{\mathfrak{X}}=(\bm{X}_{1},\ldots,\bm{X}_{n})$ that is independently drawn from $\mathcal{MN}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$ to formulate the weighted log-likelihood function $l(\bm{w},\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}}|\bm{\mathfrak{X}})$ as
@@
A proof is given in Supplement B. Based on this proposition, we obtain the matrix-variate counterpart to the multivariate MCD estimator’s objective, resulting in robust estimators of the parameters $\bm{\operatorname{M}}$, $\bm{\Sigma}^{\operatorname{row}}$, and $\bm{\Sigma}^{\operatorname{col}}$.
@@
Because the estimators in Definition (2.0.1) are MLEs based on the observations in the subset $H$, they almost surely exist and are positive definite if $h\geq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$. If $p=1$ and/or $q=1$, optimization problem (14) coincides with the optimization problem of the MCD estimator, and one obtains the univariate or multivariate MCD estimator, respectively (Rousseeuw,, 1985).

### v2

@@ Abstract
This work introduces the Matrix Minimum Covariance Determinant (MMCD) method, a novel robust location and covariance estimation procedure designed for data that are naturally represented in the form of a matrix. Unlike standard robust multivariate estimators, which would only be applicable after a vectorization of the matrix-variate samples leading to high-dimensional datasets, the MMCD estimators account for the matrix-variate data structure and consistently estimate the mean matrix, as well as the rowwise and columnwise covariance matrices in the class of matrix-variate elliptical distributions. Additionally, we show that the MMCD estimators are matrix affine equivariant and achieve a higher breakdown point than the maximal achievable one by any multivariate, affine equivariant location/covariance estimator when applied to the vectorized data. An efficient algorithm with convergence guarantees is proposed and implemented. As a result, robust Mahalanobis distances based on MMCD estimators offer a reliable tool for outlier detection. Additionally, we extend the concept of Shapley values for outlier explanation to the matrix-variate setting, enabling the decomposition of the squared Mahalanobis distances into contributions of the rows, columns, or individual cells of matrix-valued observations. Notably, both the theoretical guarantees and simulations show that the MMCD estimators outperform robust estimators based on vectorized observations, offering better computational efficiency and improved robustness. Moreover, real-world data examples demonstrate the practical relevance of the MMCD estimators and the resulting robust Shapley values.
Keywords Matrix-variate distributions  $\cdot$ Covariance with Kronecker structure  $\cdot$ Explainable artificial intelligence  $\cdot$ Image data  $\cdot$ Minimum covariance determinant  $\cdot$ Shapley values
@@ 1 Introduction
Thanks to modern data collection tools, the amount and complexity of available information are increasing rapidly, and matrix-valued data are often observed. Compared to classical multivariate observations, where values for $p$ variables are recorded for one subject, matrix-valued observations are recorded on a grid of $p\times q$ variables. These are then naturally represented as a matrix with $p$ rows and $q$ columns. Some examples include image data, where $p$ and $q$ are given by the resolution of the image, or multivariate data measured on $p$ variables, where the measurements for a subject are available for $q$ replications (e.g., different time points, different spatial locations, different experimental conditions, etc.). Frequently, matrix-valued data are analyzed as classical multivariate data by stacking the matrix columns (or rows) to a vector of length $p\cdot q$. Thus, if $n$ observations are available, the data are arranged in a matrix of dimension $n\times pq$. Depending on the dimensions, this can create high-dimensional data, possibly with a sample size lower than the resulting dimensionality, which constitutes a limitation for multivariate statistical methods.
@@
As an alternative to vectorizing matrix-valued observations, we model them under the assumption that they originate from a certain matrix-variate distribution. As in the multivariate setting, the class of matrix-elliptical distributions (Gupta and Varga,, 2012), serves as a natural ground for studying covariance estimation. The matrix-elliptical family is a semi-parametric class of distributions parametrized by the mean $\bm{\operatorname{M}}\in\mathbb{R}^{p\times q}$, row covariance $\bm{\Sigma}^{\operatorname{row}}\in\operatorname{PDS}(p)$, column covariance $\bm{\Sigma}^{\operatorname{col}}\in\operatorname{PDS}(q)$, and the so-called density generator function $g:[0,\infty)\to\mathbb{R}$. Here, $\operatorname{PDS}(a)$, with $a\in\mathbb{N}$, denotes the class of all positive definite symmetric $a\times a$ matrices. More specifically, a random matrix $\bm{X}$ with an absolutely continuous distribution has an elliptical distribution, denoted $\mathcal{ME}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}},g)$, if its density can be written as
$f(\bm{X})=\det(\bm{\Sigma}^{\operatorname{row}})^{\nicefrac{{-q}}{{2}}}\det(\bm{\Sigma}^{\operatorname{col}})^{\nicefrac{{-p}}{{2}}}g(\operatorname{tr}(\bm{\Omega}^{\operatorname{col}}(\bm{X}-\bm{\operatorname{M}})^{\prime}\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}}))),$ (1)
with $\bm{\Omega}^{\operatorname{row}}=(\bm{\Sigma}^{\operatorname{row}})^{-1}$ and $\bm{\Omega}^{\operatorname{col}}=(\bm{\Sigma}^{\operatorname{col}})^{-1}$ denoting the precision matrices among the rows and columns, respectively. Matrix elliptical distributions can also be related to their multivariate counterparts. Formally, a random matrix $\bm{X}$ follows a matrix elliptical distribution $\mathcal{ME}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}},g)$ if and only if its vectorized version $\operatorname{vec}{\bm{X}}$ follows a multivariate elliptical distribution $\mathcal{E}(\operatorname{vec}(\bm{\operatorname{M}}),\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}},g)$ (Gupta and Varga,, 2012). Here, $\operatorname{vec}(\cdot)$ is the vectorization operator, stacking the columns of a matrix on top of each other, $\otimes$ is the Kronecker product. Probably the most studied matrix elliptical distribution is the matrix normal distribution (Dawid,, 1981), denoted $\mathcal{MN}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$, with density
$f(\bm{X}|\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})=\frac{\exp(-\frac{1}{2}\operatorname{tr}(\bm{\Omega}^{\operatorname{col}}(\bm{X}-\bm{\operatorname{M}})^{\prime}\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}})))}{(2\pi)^{\nicefrac{{pq}}{{2}}}\det(\bm{\Sigma}^{\operatorname{col}})^{\nicefrac{{p}}{{2}}}\det(\bm{\Sigma}^{\operatorname{row}})^{\nicefrac{{q}}{{2}}}}.$ (2)
@@
Regarding the estimation of location and covariance for an i.i.d. sample $\bm{\mathfrak{X}}=(\bm{X}_{1},\ldots,\bm{X}_{n})\in\mathbb{R}^{n\times p\times q}$, with $\bm{X}_{i}\sim\mathcal{MN}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$, we can either work with the vectorized observations or directly with the matrices. In the former setting, the existence and uniqueness of the maximum likelihood estimator (MLE) for the covariance is guaranteed almost surely if $n\geq pq+1$. However, this approach does not take advantage of the Kronecker structure of the covariance matrix and instead directly estimates the entire $pq$-dimensional matrix $\bm{\Sigma}$. In contrast, if we utilize the knowledge of the inherent data structure, we only need to estimate the $p$-dimensional rowwise covariance matrix $\bm{\Sigma}^{\operatorname{row}}$ and the $q$-dimensional columnwise covariance matrix $\bm{\Sigma}^{\operatorname{col}}$. For the matrix-variate sample $\bm{\mathfrak{X}}$, the MLEs for the mean, as well as for the rowwise and columnwise covariance, are given by (Dutilleul,, 1999):
@@
Soloveychik and Trushin, (2016) showed that for $n$ i.i.d. samples from a continuous $p\times q$ matrix-variate distribution, there exists no unique maximum of the matrix normal likelihood function if $n<\max(\nicefrac{{p}}{{q}},\nicefrac{{q}}{{p}})+1$, and that a unique maximum exists almost surely if $n\geq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$. Although there are no closed-form solutions for the maximum likelihood estimates (MLEs) of $\bm{\Sigma}^{\operatorname{row}}$ and $\bm{\Sigma}^{\operatorname{col}}$, Dutilleul, (1999) proposed an iterative estimation procedure. The idea of the so-called flip-flop algorithm is to alternate between the computation of $\hat{\bm{\Sigma}}^{\operatorname{row}}$ and $\hat{\bm{\Sigma}}^{\operatorname{col}}$ based on Equations (4) and (5), respectively, until a convergence criterion is met. The algorithm is constructed such that positive definite estimates of subsequent iterations are nondecreasing in likelihood (Lu and Zimmerman,, 2005), and it converges almost surely to the unique maximum from any symmetric positive definite initialization of either $\hat{\bm{\Sigma}}^{\operatorname{row}}$ or $\hat{\bm{\Sigma}}^{\operatorname{col}}$, if $n\geq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$ (Soloveychik and Trushin,, 2016).
@@
Existing proposals for robust covariance estimation include a generalization of Tyler’s M-estimator (Tyler,, 1987) introduced by Soloveychik and Trushin, (2016), a robust estimator for structured covariance matrices with Kronecker structure as a particular case (Sun et al.,, 2016), distribution-free robust covariance estimation (Zhang et al.,, 2022), and ML estimation for the matrix t-distribution (Thompson et al.,, 2020).
We propose novel robust estimators for the parameters $\bm{\operatorname{M}}$, $\bm{\Sigma}^{\operatorname{row}}$, and $\bm{\Sigma}^{\operatorname{col}}$, termed the matrix minimum covariance determinant (MMCD) estimators. These estimators generalize the minimum covariance determinant (MCD) approach (Rousseeuw,, 1985), one of the most widely used approaches for robustly estimating the mean and covariance of multivariate (vector-valued) data. We show that the MMCD estimators are equivariant under matrix affine transformations and surpass the maximal attainable breakdown point of any multivariate, affine equivariant location/covariance estimator when applied to the vectorized data, such as the MCD estimator. Additionally, we show that the MMCD estimators are consistent for the finite-dimensional parameters $(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$ of the matrix elliptical distribution, thus bridging a gap between the individual, distribution-specific, estimators in the elliptical family. Furthermore, a concentration step (C-step) algorithm is developed to efficiently compute the MMCD estimators; see Rousseeuw and Driessen, (1999) for more details on C-step for MCD. Additionally, we introduce a reweighting step that preserves the properties of the MMCD estimators and greatly increases finite-sample efficiency.
@@
The robust MMCD estimators can then be employed for outlier detection using the Mahalanobis distances (Mahalanobis,, 1936) for matrix-valued observations. Because it is essential to understand the reasons for the outlyingness, we extend the concept of Shapley values introduced in Mayrhofer and Filzmoser, (2023) for outlier explanation in the multivariate case to the matrix-variate setting. Shapley values (Shapley,, 1953) are well-known from explainable AI (Lundberg and Lee,, 2017), but their computation is usually time-consuming. Our proposal is computationally efficient, and the resulting Shapley values preserve their attractive properties (Shapley,, 1953).
@@
The paper is organized as follows. In Section 2, we introduce the MMCD estimators, then proceed to derive their theoretical properties in Section 3. Section 4 is devoted to computational details for the MMCD estimators. In Section 5, we propose Shapley values for outlier explanation and present their properties. In Sections 6 and 7, we illustrate the performance of the proposed methods on numerical simulations and real-world examples. Section 8 concludes our findings. The supplementary materials contain more information on the theoretical background in this context, proofs, technical derivations, code, and additional numerical results.
@@ 2 The MMCD estimators
The MLEs given in Equations (3)-(5), much like the multivariate normal MLEs, i.e. sample mean and covariance, also serve as valid (consistent) parameter estimators in the class of elliptical distributions; see Remark 3.0.1. However, just like their multivariate counterparts, these are not robust against outlying observations. In order to obtain robust estimators for the finite-dimensional parameters $(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$ in Equation (1), we optimize the weighted version of the matrix-normal (log-)likelihood function. This principle has been similarly used in the context of other robust estimators (e.g., Neykov et al.,, 2007; García-Escudero et al.,, 2010; Kurnaz et al.,, 2018), and in particular, Raymaekers and Rousseeuw, (2023) show that the MCD estimator can be reformulated in terms of likelihood; the objective of the MCD estimator is to identify the subset of $h$ out of $n$ samples ($\nicefrac{{n}}{{2}}\leq h\leq n$) with the smallest determinant of the sample covariance matrix. This is equivalent to determining a subset of size $h$ that maximizes the multivariate normal (log-)likelihood function.
Extending the concept of the multivariate MCD approach, we introduce weights $\bm{w}=(w_{1},\dots,w_{n})\in\mathbb{R}^{n}$ for a given sample $\bm{\mathfrak{X}}=(\bm{X}_{1},\ldots,\bm{X}_{n})$ that is independently drawn from $\mathcal{MN}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$ to formulate the weighted log-likelihood function $l(\bm{w},\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}}|\bm{\mathfrak{X}})$ as
@@
A proof is given in Supplement B. Based on this proposition, we obtain a matrix-variate counterpart to the multivariate MCD estimator’s objective, resulting in robust estimators of the parameters $\bm{\operatorname{M}}$, $\bm{\Sigma}^{\operatorname{row}}$, and $\bm{\Sigma}^{\operatorname{col}}$.
@@
The estimators in Definition (2.0.1) almost surely exist and are positive definite if $h\geq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$ (Soloveychik and Trushin,, 2016). If $p=1$ and/or $q=1$, optimization problem (13) coincides with the optimization problem of the MCD estimator, and one obtains the univariate or multivariate MCD estimator, respectively.